# galaxies: active — galaxies: individual (NGC 1052) — galaxies: nuclei — ISM: absorption lines — ISM: jets and outflows Jet-torus interaction revealed by sub-parsec SO absorption in NGC 1052

### v1

@@ 3.2 SO absorption
A triple-Gaussian fit was applied to the spectrum, resulting in a per degree of freedom 36.8 / 29. The best-fit Gaussian components are overlaid on the spectrum in figure 3, and their parameters are listed in table 3.2. To assess the statistical significance of the fitted absorption components, we applied a -test under the null hypothesis that the peak amplitude () is zero. The resulting -values were less than 0.01 for all three components, confirming their statistical significant.
@@ 4.2 SO absorption line profile and optical depth
As described in section 3.2, the SO absorption in the KVN results is deeper than that of the ALMA data. The observed difference in depths may be mainly due to the different spatial resolution between the KVN ( mas, corresponding to 0.07 pc) and ALMA (, corresponding to 30 pc) observations. On sub-arcsecond scales with ALMA, the background continuum emission with the extended jets covers a region of several tens of parsecs, whereas the SO gas surrounding the SMBH is distributed within a much more compact region. Consequently, the covering factor of the gas becomes less than unity as the resolution becomes coarser.
@@
where is the optical depth distribution of absorption feature shown in figure 4, and is the 129-GHz continuum image shown in figure 1. The estimated and are listed in table 3.2, together with the SO optical depth measured with a 30 pc spatial resolution by ALMA () from figure 3 in Kameno et al. (2023a). The derived mean covering factor values are slightly lower than the covering factor yielded from the HCN absorption (0.17; Kameno et al., 2023a). However, the values of are at least three times higher than . The discrepancy in optical depth can be attributed to differences in spatial resolution. Specifically, continuum emission from the extended jet of NGC 1052 is detected with ALMA but resolved out in the KVN data. Moreover, contamination from SO emission may partially fill the absorption in the ALMA data, while the KVN spatially excludes such thermal contamination. It should be also noted that Kameno et al. (2023a) calculate the contamination-corrected SO optical depth under the assumption that the extra line equivalent width equals to the average of other SO2 lines at 128.6 and 129.5 GHz, while the SO optical depth derived from our KVN observations might be the result of possible blending with the SO2 line.
@@ 4.3 Location and kinematics of SO absorption
The absence of significant SO absorption against the outer jet components E1 and W1 places a spatial constraint on the distribution of the SO gas. The spatial extent of the SO gas must therefore be smaller than the angular separation between E1 and W1 (4.8 mas), which corresponds to an apparent size of less than 0.45 pc on the sky. The detection of absorption only toward the central source, and not toward the outer receding component W1, contrasts with the characteristics observed in the 3mm HCN and HCO+ absorptions (Sawada-Satoh et al., 2016, 2019). This concentration toward the central source is inconsistent with the SO gas being located in the molecular layer at the outer radii of the torus, where the HCN and HCO+ gases lie. The enhanced SO optical depth of the central receding component W2 suggests that the SO gas resides at or close to the inner edge ( 1 pc) of the near side of the torus. The compact distribution of the high SO optical depth seen toward W2 at multiple velocities implies that small ( 0.1 pc) individual SO gas clumps are moving at various velocities across the line of sight to W2, which lies in the immediate vicinity of the SMBH. The variations in SO optical depth across the velocity channel maps (figure 4) highlight the inhomogeneous nature of the absorbing gas, supporting a clumpy gas structure. Since the velocity of the Red1 component is close to the peak velocities of 22 GHz H2O maser emission and HCN 1–0 absorption, it is likely that Red1 arises from the ongoing infalling gas inside the torus towards the SMBH. This interpretation is in agreement with earlier results for the redshifted H2O maser and HCN absorption features (Sawada-Satoh et al., 2008, 2016). On the other hand, Blue1 and Blue2 are the signature of at least two individual outward motions in the region.
@@ 5 Conclusions
We present the first sub-pc scale images of continuum and SO absorption in the center of NGC 1052 at 129 GHz using the KVN. The continuum image reveals a bright central source and symmetric two-sided jets extending 0.75 pc, with five distinct components (E1, E2, Core, W2, W1) identified. These components are aligned along a position angle of 64∘, consistent with prior lower-frequency VLBI studies. The central source exhibits a steep spectral index ( = –1.7) between 89 and 129 GHz, indicating optically thin synchrotron emission at 129 GHz. Its brightness temperature exceeds K, further supporting a non-thermal origin. SO absorption is clearly detected toward the central components Core, W2 and marginally at E2, with no significant absorption at E1 and W1. The absorption profile spans a broad velocity range over 700 km s-1, showing three distinct velocity groups. The SO absorption is concentrated toward the central components, with a pronounced enhancement at the receding W2 component. The spatial distribution of SO absorption indicates the absorbing gas is concentrated within a region smaller than 0.45 pc, likely located close to the inner edge of the torus. The distribution and velocity structure of SO absorption suggest that the SO gas forms clumpy and small-scale ( 0.1 pc) structures, showing both inflow and outflow motions. To explain the spatial and kinematic properties of the SO absorption, we propose a multiphase torus model in which jet-torus interactions produce shock-heated regions at the inner edge of the torus, leading to SO gas evaporation and the formation of SO clumps. These clumps are entrained by the jet and transported outward, while some clumps fall back toward the SMBH, creating inflow signatures. The line of sight intersects both outflowing and infalling clumps, producing the observed absorption profile and velocity structure. Our observations also reveal that the 129 GHz SO absorption shows partial similarities to the 321 GHz H2O maser in both radial velocity and spatial distribution. The spatial coincidence between SO absorption and H2O maser emission suggests that the excitation of the H2O maser could be linked to the jet–torus interaction.

### v2

@@ 3.2 SO absorption
A triple-Gaussian fit was applied to the spectrum, resulting in a per degree of freedom 36.8 / 29. The best-fitting Gaussian components are overlaid on the spectrum in figure 3, and their parameters are listed in table 3.2. To assess the statistical significance of the fitted absorption components, we applied a -test under the null hypothesis that the peak amplitude () is zero. The resulting -values were less than 0.01 for all three components, confirming their statistical significant.
@@ 4.2 SO absorption line profile and optical depth
As described in subsection 3.2, the SO absorption in the KVN results is deeper than that of the ALMA data. The observed difference in depths may be mainly due to the different spatial resolution between the KVN ( mas, corresponding to 0.07 pc) and ALMA (, corresponding to 30 pc) observations. On sub-arcsecond scales with ALMA, the background continuum emission with the extended jets covers a region of several tens of parsecs, whereas the SO gas surrounding the SMBH is distributed within a much more compact region. Consequently, the covering factor of the gas becomes less than unity as the resolution becomes coarser.
@@
where is the optical depth distribution of absorption feature shown in figure 4, and is the 129-GHz continuum image shown in figure 1. The estimated and are listed in table 3.2, together with the SO optical depth measured with a 30 pc spatial resolution by ALMA () from figure 3 in Kameno et al. (2023a). The derived mean covering factor values are slightly lower than the covering factor yielded from the HCN absorption (0.17; Kameno et al., 2023a). However, the values of are at least three times higher than . The discrepancy in optical depth can be attributed to differences in spatial resolution. Specifically, continuum emission from the extended jet of NGC 1052 is detected with ALMA but resolved out in the KVN data. Moreover, contamination from SO emission may partially fill the absorption in the ALMA data, while the KVN spatially excludes such thermal contamination. It should also be noted that Kameno et al. (2023a) calculate the contamination-corrected SO optical depth under the assumption that the extra line equivalent width is equal to the average of other SO2 lines at 128.6 and 129.5 GHz, while the SO optical depth derived from our KVN observations might be the result of possible blending with the SO2 line.
@@ 4.3 Location and kinematics of SO absorption
The absence of significant SO absorption against the outer jet components E1 and W1 places a spatial constraint on the distribution of the SO gas. The spatial extent of the SO gas must therefore be smaller than the angular separation between E1 and W1 (4.8 mas), which corresponds to an apparent size of less than 0.45 pc on the sky. The detection of absorption only toward the central source, and not toward the outer receding component W1, contrasts with the characteristics observed in the 3mm HCN and HCO+ absorptions (Sawada-Satoh et al., 2016, 2019). This concentration toward the central source is inconsistent with the SO gas being located in the molecular layer at the outer radii of the torus, where the HCN and HCO+ gases lie. The enhanced SO optical depth of the central receding component W2 suggests that the SO gas resides at or close to the inner edge ( 1 pc) of the near side of the torus. The compact distribution of the high SO optical depth seen toward W2 at multiple velocities implies that small ( 0.1 pc) individual SO gas clumps are moving at various velocities across the line of sight to W2, which lies in the immediate vicinity of the SMBH. The variations in SO optical depth across the velocity channel maps (figure 4) highlight the inhomogeneous nature of the absorbing gas, supporting a clumpy gas structure. Since the velocity of the Red1 component is close to the peak velocities of 22 GHz H2O maser emission and HCN 1–0 absorption, it is likely that Red1 arises from the ongoing infalling gas inside the torus toward the SMBH. This interpretation is in agreement with earlier results for the redshifted H2O maser and HCN absorption features (Sawada-Satoh et al., 2008, 2016). On the other hand, Blue1 and Blue2 are the signature of at least two individual outward motions in the region.
@@ 5 Conclusions
We present the first sub-pc scale images of continuum and SO absorption in the center of NGC 1052 at 129 GHz using the KVN. The continuum image reveals a bright central source and symmetric two-sided jets extending 0.75 pc, with five distinct components (E1, E2, Core, W2, W1) identified. These components are aligned along a position angle of 64∘, consistent with prior lower-frequency VLBI studies. The central source exhibits a steep spectral index ( = –1.7) between 89 and 129 GHz, indicating optically thin synchrotron emission at 129 GHz. Its brightness temperature exceeds K, further supporting a non-thermal origin. SO absorption is clearly detected toward the central components Core, W2 and marginally at E2, with no significant absorption at E1 and W1. The absorption profile spans a broad velocity range over 700 km s-1, showing three distinct velocity groups. The SO absorption is concentrated toward the central components, with a pronounced enhancement at the receding W2 component. The spatial distribution of SO absorption indicates that the absorbing gas is concentrated within a region smaller than 0.45 pc, likely located close to the inner edge of the torus. The distribution and velocity structure of SO absorption suggest that the SO gas forms clumpy and small-scale ( 0.1 pc) structures, showing both inflow and outflow motions. To explain the spatial and kinematic properties of the SO absorption, we propose a multiphase torus model in which jet-torus interactions produce shock-heated regions at the inner edge of the torus, leading to SO gas evaporation and the formation of SO clumps. These clumps are entrained by the jet and transported outward, while some clumps fall back toward the SMBH, creating inflow signatures. The line of sight intersects both outflowing and infalling clumps, producing the observed absorption profile and velocity structure. Our observations also reveal that the 129 GHz SO absorption shows partial similarities to the 321 GHz H2O maser in both radial velocity and spatial distribution. The spatial coincidence between SO absorption and H2O maser emission suggests that the excitation of the H2O maser could be linked to the jet–torus interaction.